OpenStudy (anonymous):

Solve each equation. Check for extraneous solutions. 1)\[\sqrt[3]{x-3}=1\] 2) \[\sqrt{x+7}=x+1\] 3) \[\sqrt{3x-8}=2\] What are extraneous solutions? Solve and show steps please, this is new to me >.<

OpenStudy (bahrom7893):

square both sides in each equation and then solve

OpenStudy (amistre64):

when trying to make an equation solvable, we introduce false solutions

OpenStudy (amistre64):

the dbl chk weeds out the false solutions

OpenStudy (anonymous):

how would you square if it's the cube root

OpenStudy (amistre64):

put a ^2 to it :)

OpenStudy (amistre64):

cube each side ....

OpenStudy (anonymous):

can you do the first one and show the steps x_x

OpenStudy (amistre64):

when does cbrt(n) = 1?

OpenStudy (anonymous):

when its negative?

OpenStudy (amistre64):

lets try that out; cbrt(-8) = -2 nope

OpenStudy (amistre64):

to undo a radical you put a power to it

OpenStudy (bahrom7893):

dude come on first one is obviously 4...

OpenStudy (amistre64):

cbrt^3 undo each other sqrt^2 undo each other 6rt^6 undo each other

OpenStudy (amistre64):

whatever you do to one side, you do to the other as well

OpenStudy (anonymous):

i don't understand how to get rid of the cube root.

OpenStudy (amistre64):

\[(\sqrt[n]{\ a\ })^n=a\]

OpenStudy (amistre64):

\[(\sqrt[3]{\ x\ })^3=x\]

OpenStudy (anonymous):

ok so it would be x^3-9=3

OpenStudy (bahrom7893):

\[\sqrt[3]{x-3}=1\]\[(\sqrt[3]{x-3})^3=(1)^3\]Do you get it?

OpenStudy (amistre64):

you realy should go back a few steps in math to catch up to where you wanna be now...

OpenStudy (bahrom7893):

good idea.. so should i btw

OpenStudy (bahrom7893):

Quote: can someone help @amistre64 @myininaya @lalaly @bahrom7893 btw I'm flattered to see my name among the legends. Even though it's the last one :D!

OpenStudy (anonymous):

so the answers 1? I'm confused now

OpenStudy (anonymous):

honestly i'm just trying to get my homework done, i dont understand this though -.-

OpenStudy (bahrom7893):

I don't see how this is so hard, man seriously u need to get some real tutoring help: 1)\[\sqrt[3]{x-3}=1\]\[(\sqrt[3]{x-3})^3=(1)^3\]\[x-3=1\]\[x=1+3=4\]

OpenStudy (anonymous):

ok so for the second one squaring takes out the radical but makes the other part squared?

OpenStudy (bahrom7893):

1) \[\sqrt{x+7}=x+1\]\[(\sqrt{x+7})^2=(x+1)^2\]\[x+7=x^2+2x+1\]Now u can just simplify, factor etc..

OpenStudy (bahrom7893):

after squaring both sides for the last one u'll get: 3x-8=4 3x=12 x=4

OpenStudy (anonymous):

Can you help me with 3 more of the same type please, i should of posted these instead because they look harder

OpenStudy (anonymous):

\[(2x+1)^{1/3}=3\]

OpenStudy (anonymous):

\[\sqrt{x^2-5}=4\] \[3(x+1)^{4/3}=48\]

OpenStudy (anonymous):

-> 2x + 1 = 3^3 -> 2x = 27 - 1 => x = 26/2 = 13

OpenStudy (anonymous):

what about the 1/3 power?

OpenStudy (anonymous):

x^2 - 5 = 4^2 ->x^2 = 16 + 5 = 21 => x = +/- sqrt ( 21)

OpenStudy (anonymous):

You always power both sides!

OpenStudy (anonymous):

oh so 1/3 power is the same as being cubed ok

OpenStudy (anonymous):

3(x+1)^4/3 = 48 ->(x+1)^4/3 = 48/ 3 = 16 => x + 1 = 16 ^ (3/4) => x = 8 - 1 = 7

OpenStudy (anonymous):

Even you have the result, but you need to plug it back to check for extraneous!

OpenStudy (anonymous):

what does extraneous mean & you check for it

OpenStudy (anonymous):

how do*

OpenStudy (anonymous):

Just plug the root back to the origional see if they match, the equation is true!